The nut Case has 20 pounds of mixed peanuts and hickory nuts which sells for 8.40 per pond . Peanuts sell for 8.00 per pound, and hickory nuts sell for 9.00 per pound. How many pounds of hickory nuts are in the mix?

Answer by jorel555(1290)

Let p=peanuts. Then
8p+9(20-p)=8.40(20)
8p+180-9p=168
p=180-168=12 lbs of peanuts, 8 lbs of hickory nuts in the mix!!!!!!!!!!!!!!